# Representing Molecules with Algebraic Data Types: Beyond SMILES and SELFIES

### v4

@@ 1.2.3 Extensions
SMILES remains the dominant line notation, but fixes and extensions continue to appear. DeepSMILES reduces bracket/closure errors yet does not eliminate syntactic invalidity altogether [36]. Daylight’s ecosystem introduced related languages (e.g., SMARTS/SMIRKS for queries and transforms; CHUCKLES/CHORTLES for mixtures), and OpenSMILES provides a community specification [37]. For polymers, BigSMILES extends string notation to stochastic connectivity [16]. For reactions, SMIRKS captures structural transforms but leaves conditions to external metadata; later, RInChI added a layered, direction‑aware identifier, and ProcAuxInfo was proposed for process data (yields, temperature, concentrations) [38, 17, 39]. These ad-hoc improvements continue to this day [40]. Finally, Balsa offers a fully specified, machine‑readable subset of SMILES to reduce ambiguity across implementations [25]. These innovations improve coverage but also fragment the landscape; our focus is on representations that are precise, interoperable, and conducive to learning.
@@
Fingerprints encode features of a molecule into arrays for retrieval and modeling; we treat them here as descriptors, not as canonical molecular representations [10, 16]. Classical examples include atom pairs and topological torsions (topology), ECFP/FCFP (local neighborhoods), and pharmacophore pairs/triplets (functional interactions); string‑based variants operate directly on SMILES (e.g., LINGO; MAP4) [11, 10]. In practice, analysis relies on set/bit similarities rather than vector‑space algebra; the Tanimoto (Jaccard) index is standard [12]. Limitations are well known: tautomers may hash to different bitmaps [41], and hashing plus high dimensionality can induce bit collisions and sparsity that limit expressivity [29].
@@ 1.3.2 Functional languages and cheminformatics
We are not aware of prior work that advises cheminformaticians use constructs rooted in functional programming languages directly in the representation of molecules themselves. Additionally, unlike previous work in functional programming, we attempt to expand the scope of molecular representation, and explicitly consider the role of the data type and how this affects machine learning, including with concepts such as Bayesian machine learning and Geometric deep learning.
@@ 1.3.3 RDKit as a Representation
RDKit is a mature and widely adopted toolkit for parsing, editing, and interoperating molecular data. Our aim, however, is different: we seek a foundational representation suited to differentiable, symmetry-aware Bayesian machine learning, where invariants are explicit and transformations are compositional. Structurally, RDKit’s core abstraction is a mutable, label-based 2D molecular graph, with three-dimensional geometry stored separately in Conformer objects and stereochemistry or delocalisation encoded via enumerated labels rather than algebraic structure [46]. This separation complicates formal reasoning about global invariants (e.g., consistency between topology, geometry, and electronic interpretation) and makes it difficult to treat molecules as immutable mathematical objects. In contrast, we adopt an algebraic, typed substrate in which constitution, geometry, and electronic annotations are integrated into a single structured value, enabling deterministic traversal, principled validation, and compositional transformations. RDKit therefore plays a complementary role in our workflow—as an interoperation and import/export layer—rather than serving as the semantic core of the representation.
@@ 2.1 Our Algebraic Data Type
We implement a molecular Algebraic Data Type (ADT) in Haskell [47]—a lazy, purely (albeit potentially non‑terminating and exception‑raising) functional, type‑safe language whose algebraic data types are effectively isorecursive, exposing an implicit fold/unfold discipline via constructors and pattern matching (e.g., data List a = Nil | Cons a (List a)). Two language features are central to our design. First, type classes provide named bundles of operations that a type implementing such typeclass promises to support, i.e. instances supply the implementations; this yields principled operator overloading and generic code [48]. We use type classes to give uniform APIs for pretty‑printing, equality, hashing, and distance/selection queries over molecular parts. Second, static typing checks these contracts before execution. In the Hindley–Milner tradition, the compiler infers principal (most general) types automatically, minimizing annotation burden, while type safety (progress+preservation) ensures well‑typed programs do not “get stuck” [49, 4, 50]. We exploit this to encode chemical invariants in types and to reject ill‑formed constructions at compile time.
@@
At the core, we adopt the Dietz valence‑multigraph for molecular constitution [15]. By construction, it supports delocalised and multicentre bonding, organometallic and electron‑deficient systems, resonance alternatives, fractional/zero bond orders, ionic bonding, and explicit hydrogens. This generality addresses three limitations identified for extending SELFIES beyond organic chemistry—delocalisation, explicit hydrogen counts, and complex stereochemistry [1]. The constitution layer exposes class‑constrained constructors and pattern matches: instances capture admissible compositions (e.g., what counts as a bond or atom in a given sublanguage), while static typing rules out structurally impossible assemblies before runtime.
@@
Type‑driven composition and traversal.
Haskell type classes specify admissible molecular components and compositions, constraining construction of ill‑typed structures. Standard higher‑order and monadic combinators (e.g., map, fold, forM_) then provide idiomatic iteration, filtering, and stateful exploration over molecules, coordinates, and orbitals. This combination—type‑class abstraction plus static typing and inference—yields reusable, strongly typed building blocks for downstream algorithms without sacrificing genericity.
@@
The same typed representation extends to elementary reactions by lifting the constitution/coordinate/orbital layers over multisets of molecules with atom‑mapping annotations. Type classes define the reaction‑level interfaces (e.g., how to apply an atom map, how to aggregate stoichiometry), and static typing ensures that these lifts preserve invariants across reactants and products. To illustrate compatibility with Bayesian machine learning, we provide a reference implementation of a probabilistic model in Lazy PPL that consumes/produces values of the ADT. These implementations are architectural: they define data, interfaces, and semantics used by our experiments in later sections; evaluation results are reported separately.
@@
Beyond in‑memory computation, the ADT serves as a stable, serialisable format for storage, retrieval, and exchange of molecular information. Because molecules are constructed directly as well‑typed programs in the ADT’s grammar, learning and inference can operate on the representation itself—without ad hoc encodings or external context. Type classes deliver uniform tooling (pretty‑printing, hashing, equality), while static typing and inference deliver early error detection and refactor‑friendly guarantees that are essential for reliable cheminformatics pipelines.
@@ 2.2 Haskell
In Haskell, data defines algebraic data types (ADTs) as closed sets of constructors, so every value has a known shape and functions can pattern‑match exhaustively. Type classes declare which operations a type supports—together with intended laws that instances should satisfy—enabling ad‑hoc polymorphism (e.g., Eq for ==). Optionality is explicit via Maybe/Nothing (e.g., a missing SubShell), avoiding nulls and exception‑driven control flow. Haskell’s static types and purity shift many failures to compile time, which is valuable for scientific ML pipelines.
Our contribution is to bring these guarantees to probabilistic programming for cheminformatics. Existing Haskell Probabilistic Programming Languages (PPLs) (e.g., LazyPPL) support concise Bayesian models where lazy evaluation defers costly sampling/simulation until needed—useful for molecular property modeling and design [51, 52, 53, 54, 55, 56]. The ADT layer integrates naturally with robust graph/string encodings used in de novo design, such as SELFIES and DeepSMILES [1, 36]. (ADTs and purity are not unique to Haskell; Haskell simply offers a mature, succinct realization.)
@@ 2.3 Record Syntax
3 Applications and Evaluation
@@ 3.1 The Molecule Algebraic Data Type
Smart constructors and invariants. Updates pass through small “smart” constructors: (i) mkEdge canonicalises undirected pairs so set semantics are by construction (idempotent insertion, unambiguous lookup); (ii) mkBondingSystem takes a pool size $s\in\mathbb{N}_{0}$ and an edge set $E$, derives the cached atom scope $\mathrm{atoms}(E)$, and rejects ill‑formed inputs (negative $s$, empty $E$, or non‑canonical edges). Despite, potentially being overridden, maintaining the use of these constructors ensures the invariant $\mathrm{memberAtoms}=\mathrm{atoms}(\mathrm{memberEdges})$ and keeps systems a well‑formed finite family, in line with the paper’s requirement that bonding systems be defined over sets of atom pairs.
@@ 3.2 Example: Benzene
    Lay down the $\sigma$ framework. sigmaFramework is a Set of undirected edges (mkEdge) for the six C–C ring connections and six C–H bonds—this is the pairwise adjacency.
@@ 3.2.1 Validator
We validate every molecule immediately after construction or import, so downstream inference and learning code never explores structurally nonsensical states. The validator enforces three structural invariants—(i) every bond endpoint must refer to an existing atom, (ii) self-bonds are disallowed, and (iii) the internal bookkeeping for undirected connectivity is symmetric—and then applies a conservative, element-wise valence bound. Concretely, each atom’s “valence usage” is computed from the local $\sigma$ adjacency together with the fractional contributions implied by any Dietz bonding pools, and the molecule is rejected if any atom exceeds the element-specific maximum (getMaxBondsSymbol). Successful validation is a no-op (the moleculeis returned unchanged); failures return a short diagnostic describing which invariant was violated. The routine runs in time linear in the number of atoms and edges.
@@ 3.4 Reactions
To generalise what is written here to adequately capture the semantics of reactions and reaction networks, it would be useful to derive an implementation off of a categorical semantics such as [62]. However, this is left as future work. Chemical reaction networks align naturally with category-theoretic formalisms, where reactions can be treated as morphisms between reactant and product sets [62]. This view has been used to model biochemical pathways, open dynamical systems, and Petri nets in systems chemistry [62], making an ADT in a typed functional language with support for categorical constructs such as Functors.
@@
However, the compositional nature of the ADT, its ability to be amended to incorporate reaction properties, and its alignment with category theoretic formalisms make it a natural framework cheminformatics applications.
@@ 3.5 Storage, Transmission, and Readability
As well the computational purposes discussed elsewhere in the article, the ADT provides advantages with respect to storage of the molecule, and transmission between cheminformaticians.
@@
Parsing and construction. The SDF reader constructs the atom map from V2000 atom blocks (including coordinates and element attributes), populates localBonds from bond lines, applies formal charges from M CHG annotations, and detects common six‑membered alternating rings to build a single $\pi$ system with $s=6$ over the edges of the ring. Programmatic examples (e.g. benzene) mirror this layout—six C–C $\sigma$ edges, six C–H $\sigma$ edges, and one labelled $\pi$ system spanning the ring—so text and code paths agree on $(V,B)$.
@@ 3.6 Probabilistic Programming
This unified program couples structure generation with property supervision: the posterior concentrates mass on those atom/bond configurations that yield features consistent with the target $\log P$, while simultaneously quantifying uncertainty in the regression parameters. Practically, this provides a compact vehicle for inverse design—proposals that move the structure are immediately scored by the property model—so the chain preferentially explores chemically plausible neighborhoods that explain the target. The same grammar-and-score paradigm transfers naturally to synthesis planning: generative moves propose intermediates, and probabilistic scores (learned or mechanistic) guide search, a pattern that aligns with successful retrosynthesis strategies in the literature [73].
3.7 Existing barriers to a general molecular representations.
@@
String‑first encodings suffer from far more than localized bonding, hidden hydrogens, and patchy stereochemistry. They are syntactically brittle (SMILES can be invalid) and, even when made syntactically robust (SELFIES), remain semantically limited because they still decode to labeled 2D graphs [7, 1]. They are non‑unique and tool‑dependent (divergent canonicalisation and aromaticity perception), which injects noise into indexing and learning [16]. Their traversal‑based strings are non‑local with respect to chemistry, so string distances correlate poorly with geometric or energetic similarity [29]. Critically, they omit 3D coordinates and torsions—central for binding and stability—forcing external geometry generators and breaking round‑trips [26, 27]. Binary edges with integer orders cannot express delocalisation, multi‑centre bonding, hapticity, or zero‑order interactions (e.g., diborane, ferrocene) without ad hoc labels [17, 16, 15]. Coverage for non‑tetrahedral and axial/helical/planar stereochemistry is incomplete and inconsistent (e.g., cis/trans‑platin) and often needs 3D context to disambiguate [17, 24]. Tautomers/protomers and resonance are handled outside the representation; hydrogens are frequently implicit; ions, non‑covalent interactions, and spin are out of scope [74]. Polymers, macromolecules, crystals, and periodicity fragment into bespoke grammars; reactions lack typed support for stoichiometry, conditions, and rates (pushed into side‑channels like RInChI/ProcAuxInfo) [17, 38, 39]. Overall, strings are convenient formats, not principled data types, and they misalign with Bayesian and geometric ML, which need explicit structure, 3D symmetry actions, and typed priors/likelihoods [2, 3].
@@
The lack of quantum chemical information, which is not currently part of any modern digital representation, is also a barrier to representation, as expresed in [7]:
@@
This line of work can be read as a desire for compositional semantics and robust mutation operators. Our approach provides these by modelling molecules as typed values with explicit constructors and invariants; composition is then obtained using ordinary functions over the ADT.
@@
  • Complicated bonds (Dietz & zero-order bonds; context for FP6). We adopt the Dietz-style constitution (multigraph of bonding systems) to capture multicenter bonding and encode “zero-order” interactions as bonds with 0 shared electrons in the ADT. Krenn et al. note: “Dietz suggested a hypergraph concept …accounting for multicenter bonding,” and dashed interactions in diborane “have been termed ‘zero-order bonds’ by Clark” [1]. This work: represented as a bonding system with sharedElectrons = 0 (a “zero-order” system) and an explicit member edge set; multicentre bonding is represented by bonding systems whose member edge set has —E—¿1.

### v5

@@ 1.2.3 Extensions
SMILES remains the dominant line notation, but fixes and extensions continue to appear. DeepSMILES reduces bracket/closure errors yet does not eliminate syntactic invalidity altogether [38]. Daylight’s ecosystem introduced related languages (e.g., SMARTS/SMIRKS for queries and transforms; CHUCKLES/CHORTLES for mixtures), and OpenSMILES provides a community specification [39]. For polymers, BigSMILES extends string notation to stochastic connectivity [13]. For reactions, SMIRKS captures structural transforms but leaves conditions to external metadata; later, RInChI added a layered, direction‑aware identifier, and ProcAuxInfo was proposed for process data (yields, temperature, concentrations) [40, 15, 41]. These ad-hoc improvements continue to this day [42]. Finally, Balsa offers a fully specified, machine‑readable subset of SMILES to reduce ambiguity across implementations [27]. Extensions to SMILES/SELFIES continue to improve coverage (polymers, reactions, subsets with stricter semantics, etc.), but they also fragment the representational landscape: each extension introduces additional syntax rules and often new categorical labels, making round-trips and benchmark comparisons harder across toolchains and datasets. For ML evaluation, this matters because “validity”, augmentation strategies, and edit locality become representation- and implementation-dependent. Our focus is therefore not to propose yet another line notation, but to define a typed semantic core (MolADT) with explicit invariants and transformations, and to treat string formats as serialisations at the boundary.
@@
Fingerprints are widely used as fixed-length descriptors in classical QSAR and virtual screening benchmarks, and they highlight an important distinction for evaluation: a fingerprint is not a primary molecular representation, but a derived feature map from an underlying representation (graph/string/toolkit object). We therefore discuss fingerprints briefly to clarify where MolADT fits: MolADT is intended as the semantic substrate from which descriptors (including fingerprints) can be deterministically derived, rather than as a competing descriptor family.
Fingerprints encode molecules into fixed-length feature vectors for retrieval, clustering, and QSAR. They are effective as descriptors, but can be lossy with respect to stereochemistry and geometry and are not, by themselves, canonical molecular representations. In this paper we therefore treat fingerprints as downstream descriptors derived from a structured representation, not as the representation itself. These approaches may be effective for retrieval and similarity ranking, but they can be lossy with respect to stereochemistry or geometry, and they often require careful choices of similarity metrics and calibration when used for learning or uncertainty estimation [43, 44, 45]. In this paper we treat fingerprints as downstream descriptors rather than as a primary molecular representation. Classical examples include atom pairs and topological torsions (topology), ECFP/FCFP (local neighborhoods), and pharmacophore pairs/triplets (functional interactions); string‑based variants operate directly on SMILES (e.g., LINGO; MAP4) [44, 43]. In practice, analysis relies on set/bit similarities rather than vector‑space algebra; the Tanimoto (Jaccard) index is standard [45]. Limitations are well known: tautomers may hash to different bitmaps [46], and hashing plus high dimensionality can induce bit collisions and sparsity that limit expressivity [31].
@@ 1.3.2 Functional languages and cheminformatics
To our knowledge, most existing cheminformatics representations—regardless of implementation language—treat the molecular graph (with categorical bond labels) as the semantic core, with validation and special cases handled in library code. In contrast, MolADT is explicitly designed as a typed, layered semantic IR with first-class support for general bonding systems and well-scoped transformations; related functional-language efforts typically reimplement conventional graph models rather than rethinking the representation’s semantic contract.
@@ 1.3.3 RDKit as a Representation
RDKit is a mature and widely adopted cheminformatics toolkit that provides robust parsing, editing, and interoperability. We do not aim to replace RDKit’s algorithms. Our focus is representation-level: we propose a semantic core intended for controlled benchmarking and for workflows where explicit invariants, compositional transformations, and symmetry-aware modelling are first-class concerns.
Conceptually, RDKit’s internal model is a mutable labelled graph with 3D geometry stored separately (e.g., in conformer objects [51]), and many chemical interpretations (aromaticity, delocalisation, stereochemistry) are encoded via enumerated labels and toolkit-specific conventions [46]. This design is practical and widely successful, but it can make it harder to treat molecules as immutable mathematical objects with a single explicit set of invariants spanning constitution and geometry. MolADT instead encodes constitution, optional geometry, and optional annotations in a single typed value with explicit validation and local edits.
In this paper we use RDKit only as a point of comparison (and, where needed in future work, as an interoperability boundary); MolADT is the semantic representation under study.
@@ 2.1 Our Algebraic Data Type
We implement MolADT in Haskell, a statically typed, lazy functional programming language with native support for algebraic data types and pattern matching [52, 53, 11, 54]. This choice is pragmatic: it allows us to (i) express molecules as structured values with an explicit, checkable shape, (ii) implement validation and transformations as composable functions, and (iii) make the presence or absence of optional layers explicit to downstream code. The representation itself is language-agnostic; Haskell provides a concise reference implementation that makes the design concrete. Two language features are central to our design. First, type classes provide named bundles of operations that a type implementing such typeclass promises to support, i.e. instances supply the implementations; this yields principled operator overloading and generic code [55]. Second, static typing checks these contracts before execution.
@@
At the core, we adopt the Dietz valence‑multigraph for molecular constitution [14]. By construction, it supports delocalised and multicentre bonding, organometallic and electron‑deficient systems, resonance alternatives, fractional/zero bond orders, ionic bonding, and explicit hydrogens. This generality addresses three limitations identified for extending SELFIES beyond organic chemistry delocalisation, explicit hydrogen counts, and complex stereochemistry [1]. The constitution layer exposes class‑constrained constructors and pattern matches: instances capture admissible compositions (e.g., what counts as a bond or atom in a given sublanguage), while static typing rules out structurally impossible assemblies before runtime.
@@
Type driven composition and traversal.
Haskell type classes specify admissible molecular components and compositions, constraining construction of ill‑typed structures. Standard higher‑order and monadic combinators (e.g., map, fold, forM_) then provide idiomatic iteration, filtering, and stateful exploration over molecules, coordinates, and orbitals. This combination type class abstraction plus static typing and inference yields reusable, strongly typed building blocks for downstream algorithms without sacrificing genericity.
@@
The same typed representation extends to elementary reactions by lifting the constitution/coordinate/orbital layers over multisets of molecules with atom‑mapping annotations. Type classes define the reaction‑level interfaces (e.g., how to apply an atom map, how to aggregate stoichiometry), and static typing ensures that these lifts preserve invariants across reactants and products. To illustrate compatibility with Bayesian machine learning, we provide a reference implementation of a probabilistic model in Lazy PPL that consumes/produces values of the ADT. These implementations are architectural: they define the data, interfaces and implementations used by experiments in later sections.
@@
Beyond in‑memory computation, the ADT serves as a stable, serialisable format for storage, retrieval, and exchange of molecular information. Because molecules are constructed directly as well‑typed programs in the ADT’s grammar, learning and inference can operate on the representation itself without ad hoc encodings or external context. Type classes deliver uniform tooling (pretty‑printing, hashing, equality), while static typing and inference deliver early error detection and refactor‑friendly guarantees that are essential for reliable cheminformatics pipelines.
@@ 2.2 Haskell
In Haskell, data defines Algebraic Data Types (ADTs) as closed sets of constructors, so every value has a known shape and functions can pattern‑match exhaustively. Type classes declare which operations a type supports, together with intended laws that instances should satisfy enabling ad‑hoc polymorphism (e.g., Eq for ==). Optionality is explicit via Maybe/Nothing (e.g., a missing SubShell), avoiding nulls and exception driven control flow. Haskell’s static types and purity shift many failures to compile time, which is valuable for scientific ML pipelines.
Our contribution is to bring these guarantees to probabilistic programming for cheminformatics. Existing Haskell Probabilistic Programming Languages (PPLs) (e.g., LazyPPL) support concise Bayesian models where lazy evaluation defers costly sampling/simulation until needed useful for molecular property modeling and design [56, 57, 58, 59, 60, 61]. The ADT layer integrates naturally with robust graph/string encodings used in de novo design, such as SELFIES and DeepSMILES [1, 38]. (ADTs and purity are not unique to Haskell; Haskell simply offers a mature, succinct realization.)
@@ 2.3 Record Syntax
2.4 Use of AI-assisted editing
We used OpenAI’s GPT-5.2 Pro (ChatGPT) only for light editorial assistance (copy-editing, LaTeX hygiene, float placement). All domain content mathematics, chemistry, Haskell code, experiments, and conclusions was written by the authors; some code refactoring drew on OpenAI’s Codex. Any LLM suggestions were edited, verified, and cited. No proprietary or human-subject data were provided. The authors take full responsibility for the manuscript and codebase.
3 Reference Implementation and Worked Examples
@@ 3.1 The Molecule Algebraic Data Type
Smart constructors and invariants. Updates pass through small “smart” constructors: (i) mkEdge canonicalises undirected pairs, so set-theoretic semantics are by construction (idempotent insertion, unambiguous lookup); (ii) mkBondingSystem takes a pool size $s\in\mathbb{N}_{0}$ and an edge set $E$, derives the cached atom scope $\mathrm{atoms}(E)$, and rejects ill‑formed inputs (negative $s$, empty $E$, or non‑canonical edges). Despite, potentially being overridden, maintaining the use of these constructors ensures the invariant $\mathrm{memberAtoms}=\mathrm{atoms}(\mathrm{memberEdges})$ and keeps systems a well‑formed finite family, in line with the paper’s requirement that bonding systems be defined over sets of atom pairs.
@@ 3.2 Example: Benzene
    Lay down the $\sigma$ framework. sigmaFramework is a Set of undirected edges (mkEdge) for the six C–C ring connections and six C–H bonds this is the pairwise adjacency.
@@ 3.2.1 Validator
We validate every molecule immediately after construction or import, so downstream inference and learning code never explores structurally nonsensical states. The validator enforces three structural invariants (i) every bond endpoint must refer to an existing atom, (ii) self-bonds are disallowed, and (iii) the internal bookkeeping for undirected connectivity is symmetric and then applies a conservative, element-wise valence bound. Concretely, each atom’s “valence usage” is computed from the local $\sigma$ adjacency together with the fractional contributions implied by any Dietz bonding pools, and the molecule is rejected if any atom exceeds the element-specific maximum (getMaxBondsSymbol). Successful validation is a no-op (the molecule is returned unchanged); failures return a short diagnostic describing which invariant was violated. The routine runs in time linear in the number of atoms and edges.
@@ 3.4 Reactions
To generalise what is written here to adequately capture the syntax of reactions and reaction networks, it would be useful to derive an implementation off of a categorical semantics such as [67]. However, this is left as future work. Chemical reaction networks align naturally with category-theoretic formalisms, where reactions can be treated as morphisms between reactant and product sets [67]. This view has been used to model biochemical pathways, open dynamical systems, and Petri nets in systems chemistry [67], making an ADT in a typed functional language with support for categorical constructs such as Functors.
@@
However, the compositional nature of the ADT, its ability to be amended to incorporate reaction properties, and its alignment with category theoretic formalisms make it a natural framework for cheminformatics applications.
@@ 3.5 Storage, Transmission, and Readability
Parsing and construction. The SDF reader constructs the atom map from V2000 atom blocks (including coordinates and element attributes), populates localBonds from bond lines, applies formal charges from M CHG annotations, and detects common six‑membered alternating rings to build a single $\pi$ system with $s=6$ over the edges of the ring. Programmatic examples (e.g. benzene) mirror this layout six C–C $\sigma$ edges, six C–H $\sigma$ edges, and one labelled $\pi$ system spanning the ring so text and code paths agree on $(V,B)$.
@@ 3.6 Probabilistic Programming
This unified program couples structure generation with property supervision: the posterior concentrates mass on those atom/bond configurations that yield features consistent with the target $\log P$, while simultaneously quantifying uncertainty in the regression parameters. Practically, this provides a compact vehicle for inverse design proposals that move the structure are immediately scored by the property model so the chain preferentially explores chemically plausible neighborhoods that explain the target. The same grammar-and-score paradigm transfers naturally to synthesis planning: generative moves propose intermediates, and probabilistic scores (learned or mechanistic) guide search, a pattern that aligns with successful retrosynthesis strategies in the literature [78].
3.6.1 Why evaluation over SMILES is not sensical
In a probabilistic programming setting, the generative model and inference procedure must define a total probability distribution with a well-defined score for every execution trace. For SMILES, this forces an explicit treatment of “negative syntax”, i.e. what happens when a proposed string is not parseable as SMILES. If invalid strings are left outside the model’s domain (no likelihood defined), then standard samplers (MCMC, SMC, importance sampling) cannot correctly compute acceptance ratios or weights whenever proposals fall off-support. If, instead, invalid strings are assigned zero probability (e.g., by attempting to parse and applying a factor of -$\infty$, then nav̈e sampling over raw character sequences becomes practically unusable: proposals land in the invalid region overwhelmingly often, producing an inordinate amount of “junk”, extreme rejection rates, and highly variable weights. Either way, “sampling over SMILES strings” without an explicit syntactic failure semantics is not a fair empirical evaluation of probabilistic inference—most compute is spent rediscovering the grammar rather than exploring the molecular distribution. Consequently, principled PPL formulations must either (i) restrict the support to valid SMILES by construction (grammar-/parser-constrained generators, typed domains), or (ii) incorporate explicit parse-failure handling in the model while using proposals that respect the syntactic manifold, so that inference efficiency and reported metrics reflect modeling quality rather than accidental syntax validity.
3.7 Existing barriers to general molecular representations.
@@ 3.6.1 Why evaluation over SMILES is not sensical
Having introduced the MolADT design and its reference implementation, we now summarise the concrete representational requirements that most strongly affect ML evaluation protocols. These requirements motivate why we treat representation as a semantic contract rather than as a neutral serialisation, and they also define what should be reported (and ideally controlled) when benchmarking molecular ML models. String‑first encodings suffer from far more than localized bonding, hidden hydrogens, and patchy stereochemistry. They are syntactically brittle (SMILES can be invalid) and, even when made syntactically robust (SELFIES), remain semantically limited because they still decode to labeled 2D graphs [4, 1]. They are non‑unique and tool‑dependent (divergent canonicalisation and aromaticity perception), which injects noise into indexing and learning [13]. Their traversal‑based strings are non‑local with respect to chemistry, so string distances correlate poorly with geometric or energetic similarity [31]. Critically, they omit 3D coordinates and torsions central for binding and stability, forcing external geometry generators and breaking round‑trips [28, 29]. Binary edges with integer orders cannot express delocalisation, multi‑centre bonding, hapticity, or zero‑order interactions (e.g., diborane, ferrocene) without ad hoc labels [15, 13, 14]. Coverage for non‑tetrahedral and axial/helical/planar stereochemistry is incomplete and inconsistent (e.g., cis/trans‑platin) and often needs 3D context to disambiguate [15, 26]. Tautomers/protomers and resonance are handled outside the representation; hydrogens are frequently implicit; ions, non‑covalent interactions, and spin are out of scope [79]. Polymers, macromolecules, crystals, and periodicity fragment into bespoke grammars; reactions lack typed support for stoichiometry, conditions, and rates (pushed into side‑channels like RInChI/ProcAuxInfo) [15, 40, 41]. Overall, strings are convenient formats, not principled data types, and they misalign with Bayesian and geometric ML, which need explicit structure, 3D symmetry actions, and typed priors/likelihoods [7, 8].
@@
The lack of quantum chemical information, which is not currently part of any modern digital representation, is also a barrier to representation, as expressed in [4]:
@@
This line of work can be read as a desire for compositional syntax and semantics coupled with robust mutation operators. Our approach provides these by modelling molecules as typed values with explicit constructors and invariants; composition is then obtained using ordinary functions over the ADT.
@@
  • Complicated bonds (Dietz & zero-order bonds; context for FP6). We adopt the Dietz-style constitution (multigraph of bonding systems) to capture multicenter bonding and encode “zero-order” interactions as bonds with 0 shared electrons in the ADT. Krenn et al. note: “Dietz suggested a hypergraph concept …accounting for multicenter bonding,” and dashed interactions in diborane “have been termed ‘zero-order bonds’ by Clark” [1]. This work: represented as a bonding system with sharedElectrons = 0 (a “zero-order” system) and an explicit member edge set; multicentre bonding is represented by bonding systems whose member edge set greater than one elements.